OpenStudy (anonymous):

Solve |8 – m| + 4 = 21.

OpenStudy (anonymous):

|8 – m| + 4 = 21. |8 – m| = 21 -4 = 17 8-m = 7 -> m = 8 -7 = 1 8-m = -7 -> m = 8 + 7 = 15

OpenStudy (anonymous):

A. m = -9, m = 25 B. m = -25, m = 9 C. m = -25, m = -9 D. m = 9, m = 25

OpenStudy (anonymous):

ty!!

OpenStudy (anonymous):

sorrryyy, Thanks for reminding Luis! |8 – m| + 4 = 21. |8 – m| = 21 -4 = 17 8-m = 17 -> m = 8 -17 = -9 8 -m = -17 m = 17 + 8 = 25

Directrix (directrix):

- 9 and 25 as the others got